# 3M-Diffusion: Latent Multi-Modal Diffusion for Text-Guided Generation of Molecular Graphs

### v1

@@ 6 Conclusion
Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design. These applications call for methods that generate diverse, and ideally novel, molecules with the desired properties. We have introduced 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge. 3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions. It then reconstructs the molecular graph and atomic attributes based on the given text descriptions using the molecule decoder. 3M-Diffusion learns a probabilistic mapping from the text space to the latent molecular graph space using the latent diffusion model. The results of our extensive experiments demonstrate that 3M-Diffusion can generate high-quality, novel and diverse molecular graphs that semantically match the textual description provided. Ablation studies underscore the importance of aligning the latent representation of texts with that of molecular graphs to the performance of 3M-Diffusion. Work in progress is aimed at text-guided molecular graph generation tasks, e.g., text-guided generation of biological macromolecules, new drugs and new materials, and experimental validation of the resulting graphs through their synthesis and characterization in the lab.
@@ Appendix A Diffusion Models
$\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right)=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t},\sigma^{2}(t-1,t)=\frac{\left(1-\alpha_{t-1}\right)\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}},$ (13)
@@ Appendix B Training and Sampling Algorithm
The complete training algorithm is depicted in Algorithm 1. Firstly, the process begins by training the graph and text encoders using contrastive learning. Following that, the Variational Autoencoder is trained to facilitate the reconstruction of molecule graphs. Lastly, the Multi-modal Molecule Latent Diffusion Model is employed to train the denoising network for the sampling process. This structured approach forms the core of our training procedure. Moreover, we present the sampling method of 3M-Diffusion in Algorithm 2.
@@
2:  Initial: Graph Encoder network $E_{g}$, Text Encoder network $E_{t}$ and Decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
@@
17:  Fix Graph Encoder $E_{g}$ from Stage 1 and Text Encoder $E_{t}$ from Pretraining Stage
@@ C.2 Baselines
Then, we will also include three baselines for unconditional generation to verify the effectiveness of our model.
@@ C.3 Setup and Hyper-parameter settings
Appendix D More Visualization Results
D.1 Property Conditional Molecule Generation
In this section, we provide more prompt examples and more experiment details for Property Conditional Molecule Generation. We generate 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values, as shown in Figure 2,  6,  7 and  8. For instance, when using the prompt ”This molecule is not like a drug,” we select the top 5 molecules with the smallest QED values, as indicated. Conversely, for the prompt ”This molecule is like a drug,” we opt for the top 5 molecules with the largest QED values. This same approach is applied to solubility, where we select molecules based on logP values. Specifically, we adjust temperature setting within the MolT5 transformer model to promote the generation of diverse molecules. We carefully selected an appropriate temperature threshold that struck a balance between diversity and maintaining the validity of the generated molecules. Moreover, when dealing with MolT5, it is worth noting that it may occasionally generate multiple unconnected molecules within a single SMILES representation. In such cases, we select the molecule among these unconnected ones that best aligns with the top desired attribute. From all figures, it becomes evident that our model excels in generating molecules with improved desired properties when compared to MolT5-large.
@@ D.3 Case Study
D.4 Unconditional Molecule Generation
We further present 50 molecules sampled from the prior distribution in Figure 11. As shown in Figure 11, our sampled molecules present rich variety and structural complexity. This demonstrates our model’s ability to generate diverse, novel, and realistic molecules.

### v2

@@ 6 Conclusion
Generating molecules with desired properties is a critical task with broad applications in drug discovery and materials design. Such applications call for methods that produce diverse, and ideally novel, molecules with the desired properties. We have introduced 3M-Diffusion, a novel multi-modal molecular graph generation method, to address this challenge. 3M-Diffusion first encodes molecular graphs into a graph latent space aligned with text descriptions. It then reconstructs the molecular graph and atomic attributes based on the given text descriptions using the molecule decoder. 3M-Diffusion learns a probabilistic mapping from the text space to the latent molecular graph space using the latent diffusion model. The results of extensive experiments show that 3M-Diffusion can generate high-quality, novel and diverse molecular structures that semantically match the textual description provided. Ablation studies underscore the critical role of aligning the latent representation of text with that of molecular graphs in 3M-Diffusion.
Acknowledgments
We thank the anonymous reviewers and Prof. Dane Morgan of University of Wisconsin-Madison for suggestions that have helped us improve the paper. This work was supported in part by grants from the National Science Foundation (2020243, 2226025) to Vasant Honavar and the National Center for Advancing Translational Sciences, and the National Institutes of Health (UL1 TR002014).
Impact Statements
This paper advances machine learning approaches to text-guided generation of molecular structures, with potential application in drug design and material discovery. The resulting molecules, if successfully synthesized and experimentally characterized in the lab, could provide new drugs for combating diseases and for new materials for wide ranging applications.
@@ Appendix A Diffusion Models
$\bm{\mu}_{\theta}\left(\mathbf{z}_{t},t-1,t\right)$ $=\frac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)+\frac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t},$ (13)
$\sigma^{2}(t-1,t)$ $=\frac{\left(1-\alpha_{t-1}\right)\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}},$
@@ Appendix B Training and Sampling Algorithm
The complete training algorithm is depicted in Algorithm 1. Firstly, the process begins by training the graph and LLM encoders using contrastive learning. Following that, the Variational Autoencoder is trained to facilitate the reconstruction of molecule graphs. Lastly, the Multi-modal Molecule Latent Diffusion Model is employed to train the denoising network for the sampling process. This structured approach forms the core of our training procedure. Moreover, we present the sampling method of 3M-Diffusion in Algorithm 2.
@@
2:  Initial: Graph Encoder network $E_{g}$, LLM Encoder network $E_{t}$ and Decoder network $D$, denoising network $\hat{\mathbf{z}}_{\theta}$
@@
17:  Fix Graph Encoder $E_{g}$ from Stage 1 and LLM Encoder $E_{t}$ from Pretraining Stage
@@ C.1 Datasets Details and Statistics
Table 5: Statistics of all datasets.
Dataset #Training #Validation #Test
ChEBI-20 15,409 1,971 1,965
PubChem 6,912 571 1,162
PCDes 7,474 1,051 2,136
MoMu 7,474 1,051 4,554
@@ C.2 Baselines
Then, we will also include seven baselines for unconditional generation to verify the effectiveness of our model.
CharRNN (Segler et al., 2018): CharRNN models a distribution over the next token based on the previously generated tokens. We train this model by maximizing the log-likelihood of the training data, which is represented as SMILES strings.
VAE (Kingma & Welling, 2013): Variational Autoencoder (VAE) consists of two neural networks: an encoder and a decoder. The encoder infers a mapping from high-dimensional data to a lower-dimensional space, while the decoder maps this lower-dimensional representation back to the original high-dimensional space. It optimizes reconstruction loss and regularization term in a form of Kullback-Leibler divergence. We use SMILES as input and output representations as discussed in (Polykovskiy et al., 2020).
AAE (Makhzani et al., 2015): Adversarial Variational Autoencoder (AAE) replaces the Kullback-Leibler divergence in VAE with an adversarial objective. An auxiliary discriminator network is trained to distinguish between samples from a prior distribution and the model’s latent codes. Similar to VAE, we use SMILES as input and output representations as shown in (Polykovskiy et al., 2020).
LatentGAN (Prykhodko et al., 2019): The Latent Vector Based Generative Adversarial Network (LatentGAN) integrates an autoencoder with a generative adversarial network. It first pretrains an autoencoder to map SMILES structures to latent vectors. Subsequently, a generative adversarial network is trained to generate latent vectors for the pre-trained decoder.
@@
Note that we use the code for CharRNN, VAE, AAE, LatentGAN by the Molecular Sets (MOSES) Benchmark Polykovskiy et al. (2020).
@@ C.3 Setup and Hyper-parameter settings
Appendix D More Experimental Results
D.1 Average Inference Time Comparison
In this section, we compare the average inference times of 3M-Diffusion, MolT5, and ChemT5 by generating 1,000 samples, as shown in Figure 3. The average time to generate each molecule is calculated and presented. We observe that 3M-Diffusion achieves a speed 45 times faster than baseline methods, confirming the efficiency of our proposed model. Our proposed model not only generates molecules more quickly than previous approaches but also maintains a high level of semantic alignment with textual descriptions.
D.2 Property Conditional Molecule Generation
In this section, we provide more prompt examples and more experiment details for Property Conditional Molecule Generation. We generate 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values, as shown in Figure 2,  5,  6 and  7. For instance, when using the prompt ”This molecule is not like a drug,” we select the top 5 molecules with the smallest QED values, as indicated. Conversely, for the prompt ”This molecule is like a drug,” we opt for the top 5 molecules with the largest QED values. This same approach is applied to solubility, where we select molecules based on logP values. Specifically, we adjust temperature setting within the MolT5 transformer model to promote the generation of diverse molecules. We carefully selected an appropriate temperature threshold that struck a balance between diversity and maintaining the validity of the generated molecules. Moreover, when dealing with MolT5, it is worth noting that it may occasionally generate multiple unconnected molecules within a single SMILES representation. In such cases, we select the molecule among these unconnected ones that best aligns with the top desired attribute. From all figures, it becomes evident that our model excels in generating molecules with improved desired properties when compared to MolT5-large.
@@ D.4 Case Study
D.5 Unconditional Molecule Generation
As shown in Figure 10, our sampled molecules present rich variety and structural complexity. This demonstrates our model’s ability to generate diverse, novel, and realistic molecules.